OpenStudy (anonymous):

What statement correctly describes the key features of the graph of f(x) = 4(1/2)^(x + 1) − 3? Y-intercept of (0, −1), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, −1), starts down on the left, gets closer to y = −3 on the right Y-intercept of (0, 1), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, 1), starts down on the left, gets closer to y = −3 on the right

OpenStudy (anonymous):

@Albany_Goon

OpenStudy (welshfella):

to work out the y intercept put x = 0 into f(x) and work it out

OpenStudy (welshfella):

= 4(1/2)^(0+1) - 3 = ?

OpenStudy (anonymous):

Okay did it now how do I solve?

OpenStudy (welshfella):

what value did you get ?

OpenStudy (welshfella):

= 4(1/2)^(0+1) - 3 = 2^1 - 3

OpenStudy (anonymous):

-1

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Now what?

OpenStudy (welshfella):

so work out some values of f(x) for negative x's like x = -2 2 f(-2) = 4 (1/2)^(-2+1) - 3 = 4 / (1/2) -3 = 5 f(-3) = 4(1/2)^-2 - 3 = 4 / 1/4 - 3 = 13 so the left side of the curve will descend pretty steeply

OpenStudy (anonymous):

Oh okay so what's next?

OpenStudy (welshfella):